OpenStudy (samsan9):

The first term of an arithmetic sequence is 2 and the fourth term is 11. Find the sum of the first 50 terms.

OpenStudy (anonymous):

we first need the common difference here since 4th term is 11 so 11= 2+(4-1)d [ d =common difference of the concerned arithmetic sequence]

OpenStudy (anonymous):

can you find 'd' now sum of the first 50 terms = (50/2) *( 2*2 + (50-1)*d) now you may plug the value of d to get ur answer

OpenStudy (samsan9):

is the common difference as in like the difference from a1 to a2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

here d=3

OpenStudy (samsan9):

did you get 3775?

OpenStudy (anonymous):

correct

OpenStudy (samsan9):

thanks :D

OpenStudy (anonymous):

welcome